American Products plans to lay off 10% of its employees in its aerospace dividion and 15% of its employees in its agricultural divisions. If altogether 12% if the 3000 employees in these two divisions will be laid off, then how many employees are in each division?
Answer by ptaylor(2198)

Let x=number of employees in aerospace div
Any y=number of employees in Ag div
Now we are told that:
x+y=3000-----------------------eq1
We are also told that:
0.10x+0.15y=0.12(3000)-------------------eq2
Multiply eq1 by 0.10 and we get:
0.10x+0.10y=300 now subtract this eq from eq2 and we have
0.05y=60 divide both sides by 0.05
y=1200-------------------number of employees in the ag div
from eq1, x=3000-y; x=3000-1200;
x=1800-------------------------number of employees in the aero div
ck
0.10(1800)+0.15(1200)=0.12(3000)
180+180=360
360=360

Hope this helps----ptaylor